The 20th term of an arithmetic sequence is 44 and the common difference is 4. Find the nth term
Answer by stanbon(75887)

The 20th term of an arithmetic sequence is 44 and the common difference is 4. Find the nth term
----
a(20) = a + (19)*4 = 44
a + 76 = 44
a = -32
-----
nth term: a(n) = -32 + (n-1)4
--------------------------------
Cheers,
Stan H.